# Second- and third-order properties of multidimensional Langevin equations

### v11

@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
We consider the following Langevin system:
(1) $\dot{\mathbf{x}}=\mathbf{A}\mathbf{x}+\boldsymbol{\xi},\quad\langle\boldsymbol{\xi}(t)\rangle=\mathbf{0},\quad\langle\boldsymbol{\xi}(t)\boldsymbol{\xi}(t^{\prime})^{\mathsf{T}}\rangle=2\mathbf{D}\delta(t-t^{\prime}),$
@@
where $\boldsymbol{\xi}(t)$ is a Gaussian random term, $\mathbf{A}$ and $\mathbf{D}$ are constant, and $\delta(\cdot)$ is Dirac delta function. The covariance matrix $\mathbf{C}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ is determined from the Lyapunov equation [13]:
@@
Of interest also is the condition for detailed balance. A requirement for detailed balance is that the covariance function is symmetric, $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(\tau)^{\mathsf{T}}\rangle$. By differentiating with respect to $\tau$, we get $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$ [^4] By expanding $e^{\mathbf{A}\tau}$ in a Taylor series, it is seen that this condition implies symmetry of the covariance function.. To gain insight into this relation, we switch to “covariance-identity” coordinates in which $\mathbf{C}=\mathds{1}$ [13]. (We assume that $\mathbf{C}$ is finite and positive definite, so that such a transformation can be made.) In these coordinates, detailed balance implies that $\mathbf{A}$ is symmetric, which means it is diagonalizable by an orthogonal basis. In “covariance-identity” coordinates, an orthogonal basis means uncorrelated components (because $\langle(\mathbf{v}^{\mathsf{T}}\mathbf{x})(\mathbf{x}^{\mathsf{T}}\mathbf{w})\rangle=\mathbf{v}^{\mathsf{T}}\mathbf{w}$ for constant vectors $\mathbf{v}$, $\mathbf{w}$). By the Lyapunov equation, $\mathbf{D}$ must also be diagonal in this basis. Hence the system decomposes into a number of noninteracting one-dimensional linear systems [4]. Ref. [4] also contains good discussions about general properties of linear Gaussian systems.
@@ 2.2. Comparison of experimental data with theoretical predictions
We denote the deviation $\mathbf{M}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle_{\text{expt}}-\mathbf{C}_{\text{theo}}$, where “expt” and “theo” subscripts denote experimental and theoretical values, respectively. From now on we suppress the “theo” subscript and treat $\mathbf{C}$ as fixed. We may consider the eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$, which are real because we have a matrix (again, a second-rank mixed tensor) that is symmetric in coordinates in which $\mathbf{C}=\mathds{1}$. We may imagine an ensemble of stochastic systems, each having its own value of $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ which may differ from $\mathbf{C}$, and again denote the difference by $\mathbf{M}$. We assume that the ensemble mean of $\mathbf{M}$ is zero, and we now ask the question of what a suitable ensemble variance of $\mathbf{M}$ would be that reflects significant deviations. We may then compare the elements of the experimental value of $\mathbf{M}$ to the square root of the ensemble variance. For biological applications, a factor of 0.3 or more may be considered significant.
@@
where $\sim$ denotes ensemble expectation. We may treat $(i,j)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime})$, and multiply the l.h.s. (left-hand side) by the matrix inverse of the r.h.s. (right-hand side). We denote by $\boldsymbol{\Gamma}$ the resulting matrix:
@@
It so happens that $\boldsymbol{\Gamma}$ has at most one non-zero eigenvalue $\mu$, which is equal to its trace. (This is because in coordinates where $\mathbf{C}=\mathds{1}$, the matrix $\boldsymbol{\Gamma}$ takes the form $\mathbf{v}\mathbf{v}^{\mathsf{T}}$, and therefore has rank at most 1.) We have:
(11) $\operatorname{tr}(\boldsymbol{\Gamma})=M^{ij}M^{i^{\prime}j^{\prime}}C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}=\operatorname{tr}((\mathbf{M}\mathbf{C}^{-1})^{2}).$
@@
There is, however, a problem with Eq. (9), which is that it does not obey symmetry in the indices. We may denote by $v_{ij}$ a left eigenvector of $\boldsymbol{\Gamma}$ with eigenvalue $\mu$, which satisfies:
@@
We may assess the quantitative significance of a single element $M^{ij}$ by setting $i^{\prime}=i$, $j^{\prime}=j$ in the above, taking square roots, and comparing. For an assessment of all $M^{ij}$ collectively, we may calculate the trace of $\boldsymbol{\Gamma}$ according to this prescription:
(18) $\operatorname{tr}(\boldsymbol{\Gamma})=\mu\sim\frac{d(d+1)}{2},$
@@
Then $\mathbf{S}\mathbf{C}^{-1}$ and $\mathbf{T}\mathbf{C}^{-1}$ have purely real and purely imaginary eigenvalues, respectively. We have:
@@ 2.3. Probability current density and angular momentum
At stationarity (which is assumed throughout), the probability density is $p(\mathbf{x})\propto\exp\left({-\frac{1}{2}\mathbf{x}^{\mathsf{T}}\mathbf{C}^{-1}\mathbf{x}}\right)$ and the probability current density is $\mathbf{J}(\mathbf{x})=\boldsymbol{\Omega}\mathbf{x}p(\mathbf{x})$, where $\boldsymbol{\Omega}=\mathbf{A}+\mathbf{D}\mathbf{C}^{-1}$. The matrix $\boldsymbol{\Omega}$ is traceless and has purely imaginary eigenvalues [13] which define stochastic rotation frequencies, and the corresponding complex eigenvectors define planes in which breaking of detailed balance occurs.
The dynamics of the linear system is determined by the matrices $\mathbf{A}$ and $\mathbf{D}$. However, to characterize the system based on quantities even and odd under time-reversal, we may consider the angular momentum matrix:
@@
where the time-derivatives are interpreted in Itô sense. We also have $\mathbf{L}=2\langle\mathbf{x}\circ\dot{\mathbf{x}}^{\mathsf{T}}\rangle$ where the time-derivative is interpreted in Stratonovich sense. The system may now be characterized in terms of $\mathbf{C}$, $\mathbf{L}$, and $\mathbf{D}$.
Now we specialize to the two-dimensional case, where $\mathbf{x}=(x,y)^{\mathsf{T}}$. Here, $\boldsymbol{\Omega}$ has a single pair of purely imaginary eigenvalues ±$\pm i\omega_{\mathrm{stoch}}$. The angular momentum is $L(x,y):=\langle x\dot{y}-\dot{x}y\rangle$ (where the time-derivatives are interpreted in Itô sense). The angular momentum is related to the matrix $\boldsymbol{\Omega}$ by the relation
(30) $\mathbf{A}\mathbf{C}-\mathbf{C}\mathbf{A}^{\mathsf{T}}=\begin{pmatrix}0&-L\\ L&0\end{pmatrix}=2\boldsymbol{\Omega}\mathbf{C},$
@@
In the three-dimensional case, $\boldsymbol{\Omega}$ has at most a single pair of non-zero purely imaginary eigenvalues, and the stochastic rotation frequency is half the magnitude of the angular momentum vector evaluated in “covariance-identity” coordinates.
@@
As noted in [15], the stochastic rotation frequencies can be experimentally measured by averaging over angular motions. We can formally prove this by working in coordinates where $\mathbf{C}=\mathds{1}$ and use Stratonovich calculus to transform to polar coordinates $(r,\phi)$ [^6] If we use Itô calculus, then $\dot{\phi}$ has a term $\propto r^{-2}$, whose expectation diverges because a 2D Gaussian has $p(r)\propto r$ as $r\to 0$, and it would appear that $\langle\dot{\phi}\rangle$ does not exist. However, using Stratonovich calculus, no such problem arises.. We have:
@@
where the last equality follows from the Lyapunov equation. The entropy production rate $\dot{S}$ is related to $\mathbf{H}$ as $\dot{S}=-\operatorname{tr}(\mathbf{A}\mathbf{H})$ [13]. By a similar argument as in [13], in coordinates in which $\mathbf{D}=\mathds{1}$, we see that $\mathbf{H}$ is antisymmetric and therefore, like $\boldsymbol{\Omega}$, has purely imaginary complex conjugate eigenvalues. As an alternative derivation, following [3], we may consider that $\mathbf{D}^{-1/2}\mathbf{H}\mathbf{D}^{1/2}$ is antisymmetric and therefore $\mathbf{H}$ is similar to an antisymmetric matrix. In two dimensions the (dimensionless) eigenvalues ±$\pm ih$ are given by a simple formula similar to that of $\omega_{\mathrm{stoch}}$:
@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) in [9] to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. We explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^7] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9].:
@@
and therefore combining the terms accounted for thus far:
@@
Perhaps surprisingly, there is no $d$-dependence to leading order in $T^{-1}$ (Fig. 3). This means that it is possible for the bias to dominate the variance. The $d$-dependence in the variance presumably occurs at higher order, but we do not bother to calculate it.
@@ 3.1. Preliminaries: Definition, covariance functions, and detailed balance
(58) $\dot{\mathbf{x}}$ $=\mathbf{A}\mathbf{x}+\boldsymbol{\xi}^{\mathbf{x}},$
(59) $\dot{y}$ $=\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{x}+\xi^{y},$
(60) $\langle\boldsymbol{\xi}^{\mathbf{x}}(t)\boldsymbol{\xi}^{\mathbf{x}}(t^{\prime})^{\mathsf{T}}\rangle$ $=2\mathbf{D}^{\mathbf{x}\mathbf{x}}\delta(t-t^{\prime}),$
(61) $\langle\boldsymbol{\xi}^{\mathbf{x}}(t)\xi^{y}(t^{\prime})\rangle$ $=2\mathbf{D}^{\mathbf{x}y}\delta(t-t^{\prime}),$
@@
 [^9] As with $\mathbf{x}$, we have omitted a possible additive constant in $\dot{y}$. A system with an additive constant may be transformed into Eq. (59) by considering instead the variable $y(t)-ct$, where $c$ is a constant (in this case, $c=\langle\dot{y}\rangle$). This simplification is again used throughout. Note that such a constant must vanish when detailed balance is satisfied, since we must have in that case $\langle\dot{y}\rangle=0$.
where $\boldsymbol{\xi}^{\mathbf{x}}$ and $\xi^{y}$ are Gaussian random noise and all coefficients are constant. In this case, the angular momentum between $\mathbf{x}$ and $y$ is defined as:
@@
(64) $\mathbf{L}(\mathbf{x},y)=2\langle\mathbf{x}\dot{y}\rangle+\left\langle{\frac{\mathrm{d}[\mathbf{x},y](t)}{\mathrm{d}t}}\right\rangle=2\mathbf{C}\boldsymbol{\alpha}+2\mathbf{D}^{\mathbf{x}y}.$
@@
(65) $z:=y-\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{x},$
@@
(66) $\dot{z}=\xi^{z},\quad\langle\xi^{z}(t)\rangle=0,\quad\langle\boldsymbol{\xi}^{\mathbf{x}}(t)\xi^{z}(t^{\prime})\rangle=2\mathbf{D}^{\mathbf{x}z}\delta(t-t^{\prime}),\quad\langle\xi^{z}(t)\xi^{z}(t^{\prime})\rangle=2D^{zz}\delta(t-t^{\prime}),$
where $\xi^{z}$ is Gaussian random noise, and the transformed diffusion coefficients are:
(67) $\mathbf{D}^{\mathbf{x}z}$ $=\mathbf{D}^{\mathbf{x}y}-\mathbf{D}^{\mathbf{x}\mathbf{x}}(\mathbf{A}^{-1})^{\mathsf{T}}\boldsymbol{\alpha},$
(68) $D^{zz}$ $=D^{yy}-2\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{D}^{\mathbf{x}y}+\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{D}^{\mathbf{x}\mathbf{x}}(\mathbf{A}^{-1})^{\mathsf{T}}\boldsymbol{\alpha}.$
@@
To characterize the coupling between $\mathbf{x}$ and $y$, instead of $\boldsymbol{\alpha}$ and $\mathbf{D}^{\mathbf{x}y}$, we may prefer to use $\mathbf{L}(\mathbf{x},y)$ and $\mathbf{D}^{\mathbf{x}y}$. Here, we consider $\mathbf{L}(y,\mathbf{x}):=-\mathbf{L}(\mathbf{x},y)$ and use the same criterion as in the previous section to judge significance of angular momenta.

### v12

@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
The general time-homogeneous Itô–Langevin equation is given by:
[…]
@@
where $\mathbf{A}$ and $\mathbf{D}$ are constants. The covariance matrix $\mathbf{C}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ is determined from the Lyapunov equation [13]:
@@
Of interest also is the condition for detailed balance. A requirement for detailed balance is that the covariance function is symmetric, $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(\tau)^{\mathsf{T}}\rangle$. By differentiating with respect to $\tau$, we get $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$ [^4] By expanding $e^{\mathbf{A}\tau}$ in a Taylor series, it is seen that this condition implies symmetry of the covariance function.. To gain insight into this relation, we switch to “covariance-identity” coordinates in which $\mathbf{C}=\mathds{1}$ [13]. (We assume that $\mathbf{C}$ is finite and positive definite, so that such a transformation can be made. We make this assumption throughout, except for the case of non-stationary variables to which it is not applicable.) In these coordinates, detailed balance implies that $\mathbf{A}$ is symmetric, which means it is diagonalizable by an orthogonal basis. In “covariance-identity” coordinates, an orthogonal basis means uncorrelated components (because $\langle(\mathbf{v}^{\mathsf{T}}\mathbf{x})(\mathbf{x}^{\mathsf{T}}\mathbf{w})\rangle=\mathbf{v}^{\mathsf{T}}\mathbf{w}$ for constant vectors $\mathbf{v}$, $\mathbf{w}$). By the Lyapunov equation, $\mathbf{D}$ must also be diagonal in this basis. Hence the system decomposes into a number of noninteracting one-dimensional linear systems [4]. Ref. [4] also contains good discussions about general properties of linear Gaussian systems.
@@ 2.2. Comparison of experimental data with theoretical predictions
We denote the deviation $\mathbf{M}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle_{\text{expt}}-\mathbf{C}_{\text{theo}}$, where “expt” and “theo” subscripts denote experimental and theoretical values, respectively. From now on we suppress the “theo” subscript and treat $\mathbf{C}$ as fixed. We may consider the eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$, which are real because we have a matrix (again, a second-rank mixed tensor) that is symmetric in coordinates in which $\mathbf{C}=\mathds{1}$. We may imagine an ensemble of stochastic systems, each having its own value of $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ which may differ from $\mathbf{C}$, and again denote the difference by $\mathbf{M}$. We assume that the ensemble mean of $\mathbf{M}$ is zero, and we now ask the question of what a suitable ensemble covariance of $\mathbf{M}$ would be that reflects significant deviations. We may then compare the elements of the experimental value of $\mathbf{M}$ to the square root of the ensemble variance. For biological applications, a factor of 0.3 or more may be considered significant.
@@
where $\sim$ denotes ensemble expectation. We may treat $(i,j)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime})$, and multiply the l.h.s. (left-hand side) by the matrix inverse of the r.h.s. (right-hand side). We denote by $\bm{\Gamma}$ the resulting matrix:
@@
It so happens that $\bm{\Gamma}$ has at most one non-zero eigenvalue $\mu$, which is equal to its trace. (This is because in coordinates where $\mathbf{C}=\mathds{1}$, the matrix $\bm{\Gamma}$ takes the form $\mathbf{v}\mathbf{v}^{\mathsf{T}}$, and therefore has rank at most 1.) We have:
(12) $\operatorname{tr}(\bm{\Gamma})=M^{ij}M^{i^{\prime}j^{\prime}}C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}=\operatorname{tr}((\mathbf{M}\mathbf{C}^{-1})^{2}).$
@@
There is, however, a problem with Eq. (10), which is that it does not obey symmetry in the indices. We may denote by $v_{ij}$ a left eigenvector of $\bm{\Gamma}$ with eigenvalue $\mu$, which satisfies:
@@
We may assess the quantitative significance of a single element $M^{ij}$ by setting $i^{\prime}=i$, $j^{\prime}=j$ in the above, taking square roots, and comparing. For an assessment of all $M^{ij}$ collectively, we may calculate the trace of $\bm{\Gamma}$ according to this prescription:
(19) $\operatorname{tr}(\bm{\Gamma})=\mu\sim\frac{d(d+1)}{2},$
@@
Then $\mathbf{S}\mathbf{C}^{-1}$ and $\mathbf{T}\mathbf{C}^{-1}$ have purely real and purely imaginary eigenvalues, respectively (again by arguing using coordinates in which $\mathbf{C}=\mathds{1}$). We have:
@@ 2.3. Probability current density and angular momentum
At stationarity (which is assumed throughout to the extent applicable), the probability density is $p(\mathbf{x})\propto\exp\left({-\frac{1}{2}\mathbf{x}^{\mathsf{T}}\mathbf{C}^{-1}\mathbf{x}}\right)$ and the probability current density is $\mathbf{J}(\mathbf{x})=\bm{\Omega}\mathbf{x}p(\mathbf{x})$, where $\bm{\Omega}=\mathbf{A}+\mathbf{D}\mathbf{C}^{-1}$. The matrix $\bm{\Omega}$ is traceless and has purely imaginary eigenvalues [13] which define stochastic rotation frequencies, and the corresponding complex eigenvectors define planes in which breaking of detailed balance occurs.
The dynamics of the linear system is determined by the matrices $\mathbf{A}$ and $\mathbf{D}$. However, to alternatively characterize the system by quantities even and odd under time reversal, we may consider the angular momentum matrix:
@@
where the time-derivatives are interpreted in Itô sense. We also have $\mathbf{L}=2\langle\mathbf{x}\circ\dot{\mathbf{x}}^{\mathsf{T}}\rangle$ where the time-derivative is interpreted in Stratonovich sense (denoted by the open circle). The system may now be characterized in terms of $\mathbf{C}$, $\mathbf{L}$, and $\mathbf{D}$.
Now we specialize to the two-dimensional case, where $\mathbf{x}=(x,y)^{\mathsf{T}}$. Here, $\bm{\Omega}$ has a single pair of purely imaginary eigenvalues ±$\pm i\omega_{\mathrm{stoch}}$. The angular momentum is $L(x,y):=\langle x\dot{y}-\dot{x}y\rangle$ (where the time-derivatives are interpreted in Itô sense). The angular momentum is related to the matrix $\bm{\Omega}$ by the relation
(31) $\mathbf{A}\mathbf{C}-\mathbf{C}\mathbf{A}^{\mathsf{T}}=\begin{pmatrix}0&-L\\ L&0\end{pmatrix}=2\bm{\Omega}\mathbf{C},$
@@
In the three-dimensional case, $\bm{\Omega}$ has at most a single pair of non-zero purely imaginary eigenvalues, and the stochastic rotation frequency is half the magnitude of the angular momentum vector evaluated in “covariance-identity” coordinates.
@@
As noted in [15], the stochastic rotation frequencies can be experimentally measured by averaging over angular motions. We can formally prove this by working in coordinates where $\mathbf{C}=\mathds{1}$ (the eigenvalues of $\bm{\Omega}$ and the measured stochastic rotation frequencies are independent of the coordinate system chosen, up to a sign change) and use Stratonovich calculus to transform to polar coordinates $(r,\phi)$ [^6] If we use Itô calculus, then $\dot{\phi}$ has a term $\propto r^{-2}$, whose expectation diverges because a 2D Gaussian has $p(r)\propto r$ as $r\to 0$, and it would appear that $\langle\dot{\phi}\rangle$ does not exist. However, using Stratonovich calculus, no such problem arises.. We have:
@@
where the last equality follows from the Lyapunov equation. The entropy production rate $\dot{S}$ is related to $\mathbf{H}$ as $\dot{S}=-\operatorname{tr}(\mathbf{A}\mathbf{H})$ [13]. By a similar argument as in [13], in coordinates in which $\mathbf{D}=\mathds{1}$, we see that $\mathbf{H}$ is antisymmetric and therefore, like $\bm{\Omega}$, has purely imaginary complex conjugate eigenvalues. As an alternative derivation, following [3], we may consider that $\mathbf{D}^{-1/2}\mathbf{H}\mathbf{D}^{1/2}$ is antisymmetric and therefore $\mathbf{H}$ is similar to an antisymmetric matrix. In two dimensions the (dimensionless) eigenvalues ±$\pm ih$ are given by a simple formula similar to that of $\omega_{\mathrm{stoch}}$:
@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) in [9] to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. We explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^7] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9]. Also, we have neglected the inference of the true mean $\langle\mathbf{x}\rangle$ and have assumed this as known to be $\mathbf{0}$. This simplifies the calculations while retaining interesting effects.:
@@
and therefore combining the terms accounted for thus far (the remaining terms give higher powers of $T^{-1}$):
@@
Perhaps surprisingly, there is no $d$-dependence to leading order in $T^{-1}$ (Fig. 3). This means that it is possible for the bias to dominate the variance in high dimensions. The $d$-dependence in the variance presumably occurs at higher order in $T^{-1}$, but we do not bother to calculate it.
@@ 3.1. Preliminaries: Definition, covariance functions, and detailed balance
(59) $\dot{\mathbf{x}}$ $=\mathbf{A}\mathbf{x}+\bm{\xi}^{\mathbf{x}},$
(60) $\dot{y}$ $=\bm{\alpha}^{\mathsf{T}}\mathbf{x}+\xi^{y},$
(61) $\langle\bm{\xi}^{\mathbf{x}}(t)\bm{\xi}^{\mathbf{x}}(t^{\prime})^{\mathsf{T}}\rangle$ $=2\mathbf{D}^{\mathbf{x}\mathbf{x}}\delta(t-t^{\prime}),$
(62) $\langle\bm{\xi}^{\mathbf{x}}(t)\xi^{y}(t^{\prime})\rangle$ $=2\mathbf{D}^{\mathbf{x}y}\delta(t-t^{\prime}),$
@@
 [^9] As with $\mathbf{x}$, we have omitted a possible additive constant in $\dot{y}$. A system with an additive constant may be transformed into Eq. (60) by considering instead the quantity $y(t)-ct$, where $c$ is a constant (in this case, $c=\langle\dot{y}\rangle$). This simplification is again used throughout. Note that such a constant must vanish when detailed balance is satisfied, since we must have in that case $\langle\dot{y}\rangle=0$.
where $\bm{\xi}^{\mathbf{x}}$ and $\xi^{y}$ are zero-mean Gaussian white noise and all coefficients are constant. In this case, the angular momentum between $\mathbf{x}$ and $y$ is defined as:
@@
(65) $\mathbf{L}(\mathbf{x},y)=2\langle\mathbf{x}\dot{y}\rangle+\left\langle{\frac{\mathrm{d}[\mathbf{x},y](t)}{\mathrm{d}t}}\right\rangle=2\mathbf{C}\bm{\alpha}+2\mathbf{D}^{\mathbf{x}y}.$
@@
(66) $z:=y-\bm{\alpha}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{x},$
@@
(67) $\dot{z}=\xi^{z},\quad\langle\bm{\xi}^{\mathbf{x}}(t)\xi^{z}(t^{\prime})\rangle=2\mathbf{D}^{\mathbf{x}z}\delta(t-t^{\prime}),\quad\langle\xi^{z}(t)\xi^{z}(t^{\prime})\rangle=2D^{zz}\delta(t-t^{\prime}),$
where $\xi^{z}$ is zero-mean Gaussian white noise, and the transformed diffusion coefficients are:
(68) $\mathbf{D}^{\mathbf{x}z}$ $=\mathbf{D}^{\mathbf{x}y}-\mathbf{D}^{\mathbf{x}\mathbf{x}}(\mathbf{A}^{-1})^{\mathsf{T}}\bm{\alpha},$
(69) $D^{zz}$ $=D^{yy}-2\bm{\alpha}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{D}^{\mathbf{x}y}+\bm{\alpha}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{D}^{\mathbf{x}\mathbf{x}}(\mathbf{A}^{-1})^{\mathsf{T}}\bm{\alpha}.$
@@
To characterize the coupling between $\mathbf{x}$ and $y$, instead of $\bm{\alpha}$ and $\mathbf{D}^{\mathbf{x}y}$, we may prefer to use $\mathbf{L}(\mathbf{x},y)$ and $\mathbf{D}^{\mathbf{x}y}$. Here, we consider $\mathbf{L}(y,\mathbf{x}):=-\mathbf{L}(\mathbf{x},y)$ and use the same criterion as in the previous section to judge significance of angular momenta.